# MASQ: Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion

### v1

@@ IV-D Sim-to-real
To bridge the gap between simulation and real-world performance, we utilized domain randomization and an Actuator Network. Domain randomization involves randomizing various parameters to train a robust policy under different conditions[27, 50, 51]. These parameters include robot body mass, motor strength, joint position calibration, ground friction, restitution, orientation, and magnitude of gravity. We also independently randomize friction coefficients, including ground friction and tile height. Gravity is randomized every 8 seconds with a gravity impulse duration of 0.99 seconds. The system randomizes the time steps every 6 seconds and performs the overall randomization every 4 seconds. These measures enhance the model’s robustness and adaptability. The training process for the Actuator Network captures the non-ideal relationship between PD error and the torque realized [52], thereby improving the model’s performance in real-world applications.
@@ V Experiments
We conducted experiments using the Unitree Go2 quadrupled robot with various experiments. Section V-A introduces the experimental setup. Section V-C1 shows the experimental results of the simulation. Section V-C shows the real-world experiments and comparisons.
@@ V-A Experiments setup
In the simulation environment, we have designed two types of terrain: flat and uneven. Uneven terrain is a composite of pyramid-sloping terrain and random uniform terrain. Fifteen commands, sampled within a specified range, guide the gait-learning process. The robot learns to adapt and develop various skills by following these commands.
V-B Experiments in simulation
Figure 6: Flat Terrain
Figure 7: Uneven Terrain
Figure 8: Reward over Time (Flat Terrain)
Figure 9: Reward over Time (Uneven Terrain)
We compared the proposed learning algorithm with the PPO student-teacher baseline in simulation environments. We set up two types of simulation environments: flat and uneven terrain. We trained the quadruped robot using the curriculum learning method in both environments. The reward curves show that the proposed method achieves faster convergence and superior final performance compared to the PPO baseline. We present the experimental results as the mean of the rewards over five tests for a fair comparison.
@@ V-C Real-word experiments
This paper deploys MASQ on a quadruped robot and tests it on various terrains, including flat ground, grass, and sand, where it performs excellently. In addition, we conducted a series of robustness tests for heavy impacts and side kicks. The robot can quickly return to its normal state after being disturbed.
@@ V-C2 Robustness test
To validate the robustness of the gaits trained with our method, we conduct external disturbance tests on the robot. We test the robot in a bounding gait with a cycle period of 20ms. During robustness tests, the robot performs continuous jumps in this bounding gait. The impacts of human steps on the robot are applied to introduce disturbances during its jump cycle. Under normal conditions, the robot landed simultaneously on all four legs. We monitor this process by recording the force sensor values in the robot’s feet, thereby documenting the transition from a normal state through the induced disturbance and back to the normal state. As shown in Fig. 5, the robot returns to its normal state after experiencing a disturbance within just one gait cycle.
@@ VI Conclusion and Future Work
This paper proposes a multi-agent-based motion control system for quadruped robots, leveraging a shared-parameter actor network and a centralized critic network within the CTDE framework. Our approach, implemented in Isaac Gym, demonstrates significant improvements in training speed and robustness through curriculum learning and domain randomization, leading to efficient limb cooperation and smooth sim-to-real transitions. Experiments validate the effectiveness of the CTDE method in enhancing cooperation and efficiency in motion control for symmetric robots.
Future work will extend this approach to other symmetric robots and explore advanced domain adaptation techniques to further minimize the sim-to-real gap, enhancing performance in dynamic environments.

### v2

@@ IV-D Sim-to-real
To bridge the gap between simulation and real-world performance, we used domain randomization and an Actuator Network[52]. Domain randomization involves randomizing various parameters to train a robust policy under different conditions[30, 54, 55]. These parameters include robot body mass, motor strength, joint position calibration, ground friction, restitution, orientation, and magnitude of gravity. We also independently randomize friction coefficients like ground friction. Gravity is randomized every 8 seconds with a gravity impulse duration of 0.99 seconds. Time steps are randomized every 6 seconds, with the overall randomization occurring every 4 seconds. These measures enhance the model’s robustness and adaptability. The training process for the Actuator Network captures the non-ideal relationship between PD error and the torque realized [52], thereby improving the model’s performance in real-world applications.
@@ V Experiments
We conducted experiments using the Unitree Go2 quadrupled robot with various experiments. Section V-A proposes the experimental setup. Section V-B shows the experimental results of the simulation. Section V-C shows the real-world experiments and comparisons.
Figure 5: Various terrains
@@ V-A Experiments Setup
In the simulation environment, we designed two types of terrain: flat and uneven, shown in Fig. 5. Uneven terrain is a composite of pyramid-sloping terrain and random uniform terrain. Fifteen commands, such as linear velocity, angular velocity and height of the base, sampled within a specified range, guide the gait-learning process. The robot learns to adapt and develop various skills by following these commands.
V-B Simulation Experiments
Figure 6: Return on various terrains
Figure 7: Contact force of four legs in different gaits
The MASQ was compared with the PPO student-teacher baseline in simulation environments, including two types: flat and uneven terrain. The quadruped robot was trained using the curriculum learning method, with hyperparameters such as the learning rate tuned for optimal performance in both scenarios. Reward curves in Fig. 6 show that the proposed method achieves faster convergence and better final performance compared to the PPO baseline. Experimental results are presented as the mean of rewards over five tests for a fair comparison.
@@ V-C Real-word Experiments
This paper deploys MASQ on a quadruped robot and tests it on various terrains, including flat ground, grass, and sand in Fig. 1, where it performs excellently. In addition, we conducted a series of robustness tests for heavy impacts and side kicks. The robot can quickly return to its normal state after being disturbed.
@@ V-C2 Robustness test
To validate the robustness of the gaits trained with our method, we conduct external disturbance tests on the robot. We test the robot in a bounding gait with a cycle period of 20ms. During robustness tests, the robot performs continuous jumps in this bounding gait. The impacts of human steps on the robot are applied to propose disturbances during its jump cycle. Under normal conditions, the robot landed simultaneously on all four legs. We monitored this process by recording the force sensor values in the robot’s feet, thereby documenting the transition from a normal state through the induced disturbance and back to the normal state. As shown in Fig. 4, the robot returns to its normal state after experiencing a disturbance within just one gait cycle.
@@ VI Conclusion and Future Work
This paper proposed a multi-agent-based motion control system for quadruped robots, utilizing a shared-parameter actor network and a centralized critic network within the CTDE framework. The proposed approach, implemented in Isaac Gym, demonstrated substantial improvements in training speed, robustness, and final performance, benefiting from curriculum learning and domain randomization. These advances enabled efficient limb coordination and smoother sim-to-real transitions. Experimental results confirmed the effectiveness of the method in enhancing both performance and efficiency in motion control for symmetric robots. Future work will extend the approach to other symmetric robots and explore its application in more complex dynamic environments.